OpenStudy (anonymous):

Suppose R = 10, C = .02, Q(0) = 0, E(t) = 100e^-5t. Find Q(t) and I(t).

OpenStudy (anonymous):

R(dQ/dt) + (1/C)*Q = E(t) 10Q' + 50Q = 100e^-5t

OpenStudy (anonymous):

10(Q' + 5Q) = 100e^-5t -> Q' + 5Q = 10e^-5t

OpenStudy (amistre64):

solve the homogenous first

OpenStudy (amistre64):

10Q' + 50Q = 100e^-5t Q' + 5Q = 10e^-5t Q' + 5Q = 0 Q' =-5Q Q'/Q = -5 ln(Q) = -5t + C Q = Ce^-5t

OpenStudy (amistre64):

then solve the particular

OpenStudy (amistre64):

Q = (C + 10t) e^-5t Q(0) = 0 0 = (C + 0) 1 Q = 10t e^-5t maybe .....

OpenStudy (anonymous):

Wait. Why is the particular Q = (C + 10t)e^-5t?